### v1

@@ 4 Neural NOD for Interactive Robotics
4.1 Automatic Tuning of Game Costs using NOD
Our key insight into opinion states is that they not only model agent $i$’s attitude towards option $\ell$ (determined by the sign of ${z}^{i}_{\ell}$) but also indicate how strongly the agent prefers such an option (determined by the magnitude of ${z}^{i}_{\ell}$). Therefore, we propose to use an opinion state ${z}^{i}_{\ell}$ in a dynamic game as the coefficient of the cost term that corresponds to option $\ell$. Consequently, we may use NOD to evolve the opinion state, thus automatically tuning the game cost parameter $\theta^{i}$ in (2) such that each agent can rapidly adapt to other agents’ opinion shift as well as changes in the environment, yielding situational awareness, a crucial property for interactive robotics. To this end, we define a class of opinionated game stage costs in (2), which encodes agents’ preferences through their opinion states:
$c_{t}^{i}({x}_{t},{u}_{t};{z}^{i}_{t}):=\bar{c}_{t}^{i}({x}_{t},{u}_{t})+\sum_{\ell\in\mathcal{I}_{o^{i}}}\underbrace{{z}^{i}_{\ell,t}}_{=\theta^{i}_{\ell,t}}c_{\ell,t}^{i}({x}_{t},{u}_{t}),$ (6)
@@ 4.1 Automatic Tuning of Game Costs using NOD
We illustrate our technical approach with an autonomous racing example. We conduct simulated races on the 1:1 reconstructed Thunderhill Raceway located in Willows, CA, USA (see Fig. 6). In the race, both ego and rival vehicles are constrained to remain within the track boundaries (i.e., at least one wheel is inside the track limit), while only the ego vehicle has the responsibility to avoid a collision. We model the vehicle motion using the 4D kinematic bicycle model [54]. Similar to [47], ego’s basis cost terms (and associated weights) include incentivizing overtaking ($\theta^{e}_{\rm{ov}}$), following ($\theta^{e}_{\rm{fl}}$), deviating to the inside ($\theta^{e}_{\rm{in}}$) or outside ($\theta^{e}_{\rm{ot}}$); and the rival ones include incentivizing blocking the ego ($\theta^{r}_{\rm{bl}}$), and deviating to the inside ($\theta^{r}_{\rm{in}}$) or outside ($\theta^{r}_{\rm{ot}}$). This leads to a 7D joint opinion space. The residual cost terms of each player capture optimizing lap time, enforcing safety, and regulating control efforts.
4.2 Neural Synthesis of NOD from Inverse Dynamic Games
The key question of using NOD in an ODG (6) for physical interaction is: How should NOD parameters change adaptively in response to the evolution of physical state ${x}$? Among the few attempts to address this issue, recent work [22] captures physical information in NOD parameter design via value functions of a dynamic game, assuming that each agent faces a set of mutually exclusive options. Cathcart et al. [9] uses a manually-designed NOD model for human–robot corridor passing. In this section, we present an inverse game approach to synthesize a Neural NOD model for general ODG s with no additional assumptions on each agent’s specific task and set of options.
A discrete-time Neural NOD model (c.f. (5)) is defined as:
@@ 4.2 Neural Synthesis of NOD from Inverse Dynamic Games
where $\eta_{t}=(\eta^{1}_{t},\ldots,\eta^{N_{a}}_{t},{\lambda}_{t})$, with $\eta^{i}_{t}=({d}^{{i}}_{t}$, ${b}^{i}_{t}$, $\alpha^{{i}}_{t}$, $\gamma^{{i}}_{t}$, $\beta^{{i}}_{t}$, $\delta^{i}_{t})$, is the vector that aggregates all NOD parameters. In our proposed scheme, $\eta_{t}$ is predicted by DNN $h_{\eta}$ with parameters $\phi$ and input ${x}_{t}$. Given an initial state ${x}_{0}$, the NOD model is initialized with opinion ${z}_{0}$, which is predicted by a separate DNN $h_{{z}_{0}}$ parameterized by $\phi_{0}$. The predicted ${z}_{0}$ can also be viewed as prior information of agents’ opinions, with which we can initialize NOD with a more informative opinion than a neutral one (i.e., ${z}_{0}=0$).
@@
The Neural NOD framework is agnostic to the specific DNN feature space or architecture. For ease of exposition, we use the current state ${x}_{t}$ as the input feature of DNN $h_{\eta}(\cdot;\phi)$. Alternatively, we may use an LSTM [17] or transformer [45] model for $h_{\eta}(\cdot;\phi)$, which would then predict NOD parameters based on a state history.
Our technical approach towards synthesizing a Neural NOD model is inspired by recent work on inverse dynamic games, but incorporates several key distinctions. Existing inverse game approaches, such as [42, 29, 33], are predominantly limited to learning a static cost parameter $\theta$ from a single gameplay trajectory data. As a result, the inverse game problem needs to be re-solved when the environment (e.g., the initial condition or the other agent’s intent) changes, presenting computational hurdles for real-time deployment. In contrast, our proposed approach involves offline training of a Neural NOD model from a dataset containing many gameplay trajectories. Subsequently, this model is deployed online in an ODG to dynamically adjust game cost weights based on evolving physical states, thereby facilitating situational awareness among the gameplay agents. The inverse game training needs not be performed online, provided that the deployment scenario falls within the generalization capability of the DNN that parameterizes the NOD. In addition, the online computation speed of agents’ policies via solving an ODG with learned Neural NOD is close to that of a standard dynamic game, as long as the DNN used by Neural NOD adopts fast inference. The Neural NOD framework also differs from the game-induced NOD (GiNOD) [22] in that GiNOD assumes mutually exclusive options, requires (potentially expensive) online optimization to synthesize NOD parameters, and is incapable of leveraging prior knowledge such as expert demonstrations. Our approach overcomes these drawbacks by learning a generic DNN-parameterized NOD model from interaction data.
@@
where $L$ is the imitation objective defined by a set of (partial) observations $\{\mathbf{{y}}_{n}\}_{n\in[N]}$ and the corresponding set of state and control trajectories $\{({\mathbf{x}}_{n},{\mathbf{u}}_{n})\}_{n\in[N]}$, wherein each pair is at an equilibrium of the ODG governed by Neural NOD with parameter $(\phi,\phi_{0})$. These parameters can be learned by solving MLE (8) using standard gradient-based methods [8, 27], in which gradients $\nabla_{\phi}L$ and $\nabla_{\phi_{0}}L$ can be obtained with automatic differentiation (e.g., [40, 34]) by backpropagating through the inverse game computation graph, as shown in Figure 2. Note that MLE (8) learns DNN $h_{\eta}$ and $h_{{z}_{0}}$, which take as input arbitrary state ${x}\in\mathbb{R}^{n_{x}}$, while MLE (3) only works for a specific initial state ${x}_{0}$. Our proposed training pipeline solely requires that the ODG is formulated with any existing differentiable dynamic game frameworks, such as open-loop Nash [41], feedback Nash [29], generalized Nash [33], stochastic Nash [36, 32], open-loop Stackelberg [19], and feedback Stackelberg [30].
@@
We use inverse game approach [29] with a negative log-likelihood objective, which yields an approximate feedback Nash equilibrium for the ODG. We parameterize the Neural NOD as a full-connected feed-forward DNN, of which the features include all vehicles’ physical states. We implement the training pipeline with Flax [16] and train the Neural NOD with Adam [27].
4.3 Properties of Neural NOD
In this section, we provide stability analysis of Neural NOD around the neutral opinion ${z}=0$, i.e., where all agents are indecisive. In particular, we show in the following theorem that the learned self-reinforcement gains $\alpha$ and attention ${\lambda}$ can be adjusted analytically so that ${z}=0$ is an unstable equilibrium of Neural NOD, that is, indecision (i.e., deadlocks) can be broken with an arbitrarily small input from the bias ${b}$. We start with a technical lemma, which ensures that pitchfork bifurcation is possible with a Neural NOD model.
@@ 4.3 Properties of Neural NOD
Define Jacobian matrix ${\mathbf{J}}_{0}=\left.{\mathbf{J}}(\operatorname{col}(\{{S}^{i}({z}^{{i}};\eta^{i})\}_{{i}\in\mathcal{I}_{a}}))\right|_{z=0}$. If $\exists{i}\in\mathcal{I}_{a}$, $\ell\in\mathcal{I}_{o^{i}}$ such that $\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})>0$, where $\bar{{\mathbf{J}}}_{0}=\left.{\mathbf{J}}(\operatorname{col}(\{\bar{{S}}^{i}({z}^{{i}};\eta^{i})\}_{{i}\in\mathcal{I}_{a}}))\right|_{z=0}$, $\bar{{S}}^{i}({z}^{{i}};\eta^{i})={S}_{1}\Big{(}\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\gamma^{{i}{j}}_{\ell}{z}^{{j}}_{{\ell}}\Big{)}+\sum_{{p}\in\mathcal{I}_{\theta_{i}}\setminus\{{\ell}\}}{S}_{2}\Big{(}\beta_{\ell{p}}^{{i}}{z}^{{i}}_{{p}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\delta^{{i}{j}}_{\ell{p}}{z}^{{j}}_{{p}}\Big{)}$, and $\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})$ is the eigenvalue of $\bar{{\mathbf{J}}}_{0}$ at the same location in the spectral matrix of $\bar{{\mathbf{J}}}_{0}$ as $\alpha^{i}_{\ell}$ in $\mathcal{A}=\operatorname{diag}\{\alpha^{i}_{\ell}\}_{i\in\mathcal{I}_{a},\ell\in\mathcal{I}_{o^{i}}}$, then ${\mathbf{J}}_{0}$ has at least one eigenvalue with a positive real part, i.e., ${}^{+}(\sigma({\mathbf{J}}_{0}))\neq\emptyset$.
Proof
Note that ${\mathbf{J}}_{0}=\mathcal{A}+\bar{{\mathbf{J}}}_{0}$. Since $\mathcal{A}$ is diagonal, each eigenvalue of ${\mathbf{J}}_{0}$ takes the form of $\sigma^{i}_{\ell}({{\mathbf{J}}}_{0})=\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})$. Therefore, by choosing one $\alpha^{i}_{\ell}$ such that $\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})>0$, we can ensure that ${\mathbf{J}}_{0}$ has at least one eigenvalue with positive real part. ∎
Now, we are ready to state our main theoretical result, which guarantees breaking of indecisions (deadlocks) when using a Neural NOD model for game-theoretic planning.
Theorem 4.1 (Guaranteed Indecision Breaking)
@@
  • When ${b}=0$, if ${\lambda}>\|\operatorname{col}(\{{d}_{i}\}_{{i}\in\mathcal{I}_{a}})\|_{\infty}/\max{}^{+}(\sigma({\mathbf{J}}_{0}))$, then the neutral opinion ${z}=0$ is an unstable equilibrium of the Neural NOD. Moreover, the opinion state departs ${z}=0$ (locally) at an exponential rate with an arbitrarily small bias ${b}$,
  • When ${b}\neq 0$, the neutral opinion ${z}=0$ is not an equilibrium of the Neural NOD and the pitchfork unfolds, i.e., the model is ultrasensitive at ${z}=0$.
Proof
Linearizing Neural NOD at neutral opinion ${z}=0$ gives $\dot{{z}}=(-D+{\mathbf{J}}_{0}){z}+{b}$, where $D=\operatorname{blkdiag}(\{D^{i}\}_{i\in\mathcal{I}_{a}})$. When ${b}=0$, indecision ${z}=0$ is an equilibrium. Since ${}^{+}(\sigma({\mathbf{J}}_{0}))\neq\emptyset$, if the attention variable ${\lambda}$ is chosen such that ${\lambda}>{\lambda}^{*}=\|\operatorname{col}(\{{d}_{i}\}_{{i}\in\mathcal{I}_{a}})\|_{\infty}/\max{}^{+}(\sigma({\mathbf{J}}_{0}))$, then matrix $-D+{\mathbf{J}}_{0}$ has one eigenvalue with positive real part, i.e., indecision ${z}=0$ is exponentially unstable. Thus, an indecision-breaking pitchfork bifurcation occurs at critical attention value ${\lambda}^{*}$. When ${b}\neq 0$, the pitchfork bifurcation unfolds, as predicted by the unfolding theory [15]. ∎
4.4 Using Neural NOD for Split-Second Decision-Making
Once a Neural NOD model is learned in an offline manner from data, it can be deployed in a plug-and-play fashion in an online ODG solver for fast and flexible multi-agent decision-making Moreover, Theorem 4.1 offers a systematic approach to avoid indecisions in an ODG by modifying the self-reinforcement gains $\alpha$ and attention ${\lambda}$ as a function of Neural NOD parameters. Finally, our proposed game-theoretic decision-making framework is flexible to mix decisive and indecisive players. For example, in car racing, we would like to learn a policy for a decisive ego agent, while not assuming the rivals are equally decisive at all times. In this case, we can jointly learn an opinion dynamics model for the ego and rivals, where the ego uses a Neural NOD model for fast and decisive motion planning, and rivals use a general DNN $\tilde{g}$ to generate the next opinion state based on the current opinion and physical state, i.e., ${z}^{r}_{t+1}=\tilde{g}({z}_{t},{x}_{t})$. By removing the inductive bias of NOD fast decision-making for the rival agents, we can more effectively predict their actions when they are slow at making decisions.
5 Simulation Results
We use our proposed Neural NOD model, trained on both synthetic and human datasets, to automatically tune the game cost weights within the ILQGame planning framework [14], which computes an approximate feedback Nash equilibrium strategy [5, Ch. 6] for each agent. We evaluate its performance in simulated autonomous racing scenarios (running example). For both planning and simulation, we employ the 4D kinematic bicycle model [54], discretized with a time step of $\Delta t=0.1$ s, to describe the motion of the vehicles. All planners are implemented using JAX [34] and run in real-time at a frequency of 10Hz on a desktop with an AMD Ryzen 9 7950X CPU.
Hypotheses. We make three hypotheses that highlight the advantages of Neural NOD.
  • H1 (Performance). The Neural NOD model leads to safer and more efficient robot motion than non-NOD baselines.
  • H2 (Generalization). The Neural NOD model generalizes better to other agents’ behaviors that are out of distribution.

### v2

@@ III Neural NOD for Interactive Robotics
III-A Automatic Tuning of Game Costs using NOD
Our key insight into opinion states is that they not only model agent $i$’s attitude towards option $\ell$ (determined by the sign of ${z}^{i}_{\ell}$) but also indicate how strongly the agent prefers such an option (determined by the magnitude of ${z}^{i}_{\ell}$). Therefore, we propose to use an opinion state ${z}^{i}_{\ell}$ in a dynamic game as the coefficient of the cost term that corresponds to option $\ell$. Consequently, we may use NOD to evolve the opinion state, automatically tuning cost parameter $\theta^{i}$ in (2) such that each agent can rapidly adapt to changes in the environment, a crucial property for interactive robotics. To this end, we define a class of opinionated stage costs in (2), encoding agents’ preferences through their opinion states:
$c_{t}^{i}({x}_{t},{u}_{t};{z}^{i}_{t}):=\bar{c}_{t}^{i}({x}_{t},{u}_{t})+\textstyle\sum_{\ell\in\mathcal{I}_{o^{i}}}{z}^{i}_{\ell,t}c_{\ell,t}^{i}({x}_{t},{u}_{t}),$ (6)
@@ III-A Automatic Tuning of Game Costs using NOD
We illustrate our technical approach with a simulated racing example on the 1:1 reconstructed Thunderhill Raceway located in Willows, CA, USA (see Fig. 5). In the race, both ego and rival vehicles are constrained to remain within the track boundaries (i.e., at least one wheel is inside the track limit), while only the ego vehicle has the responsibility to avoid a collision. We model the vehicle motion using the 4D kinematic bicycle model [zhang2020optimization]. Similar to [song2021autonomous], the ego’s basis cost terms (and associated weights) include incentivizing overtaking ($\theta^{e}_{\rm{ov}}$), following ($\theta^{e}_{\rm{fl}}$), deviating to the inside ($\theta^{e}_{\rm{in}}$) or outside ($\theta^{e}_{\rm{ot}}$); and the rival ones include incentivizing blocking the ego ($\theta^{r}_{\rm{bl}}$), and deviating to the inside ($\theta^{r}_{\rm{in}}$) or outside ($\theta^{r}_{\rm{ot}}$). The residual cost terms of each player capture optimizing lap time, enforcing safety, and regulating control efforts.
III-B Neural Synthesis of NOD from Inverse Dynamic Games
The key question of using NOD in an ODG (6) for physical interaction is: How should NOD parameters change adaptively in response to the evolution of physical state ${x}$? Our central contribution is an inverse game approach to synthesize a Neural NOD model for general ODG s with no additional assumptions on each agent’s specific task and set of options. Building on discrete-time NOD model (5), we may define a Neural NOD model as:
@@ III-B Neural Synthesis of NOD from Inverse Dynamic Games
where $\eta_{t}=(\eta^{1}_{t},\ldots,\eta^{N_{a}}_{t},{\lambda}_{t})$, with $\eta^{i}_{t}=({d}^{{i}}_{t}$, ${b}^{i}_{t}$, $\alpha^{{i}}_{t}$, $\gamma^{{i}}_{t}$, $\beta^{{i}}_{t}$, $\delta^{i}_{t})$, is the vector that aggregates all NOD parameters. In our proposed scheme, $\eta_{t}$ is predicted by DNN $h_{\eta}$ with parameters $\phi$ and input ${x}_{t}$. Given an initial state ${x}_{0}$, the NOD model is initialized with opinion ${z}_{0}$, which is predicted by a separate DNN $h_{{z}_{0}}$ parameterized by $\phi_{0}$. The predicted initial opinion state ${z}_{0}$ can be interpreted as the prior information of agents’ opinions, which we can use to initialize the NOD model with a more informative opinion than a neutral one (i.e., ${z}_{0}=0$) commonly used in the literature.
@@
The Neural NOD framework is agnostic to the specific DNN feature space or architecture. For ease of exposition, we use the current state ${x}_{t}$ as the input feature of DNN $h_{\eta}(\cdot;\phi)$. Alternatively, we may use an LSTM [hochreiter1997long] or transformer [shi2022mtr] model for $h_{\eta}(\cdot;\phi)$, which would then predict NOD parameters based on a state history.
Our technical approach towards synthesizing a Neural NOD model is inspired by recent work on inverse dynamic games, but incorporates several key distinctions. Existing inverse game approaches, such as [peters2023online, li2023cost, liu2023learning], are predominantly limited to learning a static cost parameter $\theta$ from a single gameplay trajectory data. As a result, the inverse game problem needs to be solved repeatedly when the environment (e.g., the initial condition or the other agent’s intent) changes, presenting computational hurdles for real-time deployment. In contrast, our proposed approach involves offline training of a Neural NOD model from a dataset containing batches of gameplay trajectories. The online computation speed of agents’ policies via solving an ODG with learned Neural NOD is close to that of a standard dynamic game, as long as the DNN used by Neural NOD adopts fast inference.
@@
where $L$ is the imitation objective defined by a set of observations $\{\mathbf{{y}}_{n}\}_{n\in[N]}$ and the corresponding set of state and control trajectories $\{({\mathbf{x}}_{n},{\mathbf{u}}_{n})\}_{n\in[N]}$, wherein each pair is at an equilibrium of the ODG governed by Neural NOD with parameter $(\phi,\phi_{0})$. These parameters can be learned by solving MLE (8) using standard gradient-based methods [bottou2010large, kingma2014adam], in which gradients $\nabla_{\phi}L$ and $\nabla_{\phi_{0}}L$ can be obtained with automatic differentiation (e.g., [paszke2017automatic, jax2018github]) by backpropagating through the inverse game computation graph, as shown in Figure 2. Note that MLE (8) learns DNN $h_{\eta}$ and $h_{{z}_{0}}$, which take as input arbitrary state ${x}\in\mathbb{R}^{n_{x}}$, while MLE (3) only works for a specific initial state ${x}_{0}$. Our proposed training pipeline solely requires that the ODG is formulated with any existing differentiable dynamic game frameworks, such as open-loop Nash [peters2020inference], feedback Nash [li2023cost], generalized Nash [liu2023learning], stochastic Nash [mehr2023maximum, lidard2024blending], open-loop Stackelberg [hu2024plays], and feedback Stackelberg [li2024computation]. Once a Neural NOD model is learned offline from data, it can be deployed plug-and-play in an online ODG solver for time-sensitive multi-agent decision-making.
@@
We use the inverse game approach [li2023cost] with a negative log-likelihood objective, which yields an approximate feedback Nash equilibrium for the ODG.
III-C Properties of Neural NOD
In this section, we show that the self-reinforcement gains $\alpha$ and attention ${\lambda}$ can be adjusted analytically so that ${z}=0$ is an unstable equilibrium of Neural NOD, that is, indecision (i.e., deadlocks) can be broken with an arbitrarily small bias term ${b}$. We start with a technical lemma, which ensures that a pitchfork bifurcation is possible with a Neural NOD model.
@@ III-C Properties of Neural NOD
If there exist ${i}\in\mathcal{I}_{a}$ and $\ell\in\mathcal{I}_{o^{i}}$ such that $\alpha^{i}_{\ell}+\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})>0$, where Jacobian matrix $\bar{{\mathbf{J}}}_{0}$ is defined as $\left.{\mathbf{J}}(\operatorname{col}(\{\bar{{S}}^{i}({z}^{{i}};\eta^{i})\}_{{i}\in\mathcal{I}_{a}}))\right|_{z=0}$, $\bar{{S}}^{i}({z}^{{i}};\eta^{i}):={S}_{1}\Big{(}\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\gamma^{{i}{j}}_{\ell}{z}^{{j}}_{{\ell}}\Big{)}+\sum_{{p}\in\mathcal{I}_{\theta_{i}}\setminus\{{\ell}\}}{S}_{2}\Big{(}\beta_{\ell{p}}^{{i}}{z}^{{i}}_{{p}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}$ $\delta^{{i}{j}}_{\ell{p}}{z}^{{j}}_{{p}}\Big{)}$, and $\sigma^{i}_{\ell}(\bar{{\mathbf{J}}}_{0})$ is the eigenvalue of $\bar{{\mathbf{J}}}_{0}$ at the same location in the spectral matrix of $\bar{{\mathbf{J}}}_{0}$ as $\alpha^{i}_{\ell}$ in diagonal matrix $\mathcal{A}=\operatorname{diag}\{\alpha^{i}_{\ell}\}_{i\in\mathcal{I}_{a},\ell\in\mathcal{I}_{o^{i}}}$, then Jacobian matrix ${\mathbf{J}}_{0}:=\left.{\mathbf{J}}(\operatorname{col}(\{{S}^{i}({z}^{{i}};\eta^{i})\}_{{i}\in\mathcal{I}_{a}}))\right|_{z=0}$ has at least one eigenvalue with a positive real part, i.e., ${}^{+}(\sigma({\mathbf{J}}_{0}))\neq\emptyset$.
Theorem 1 (Guaranteed Indecision Breaking)
@@
  • When bias term ${b}=0$, if ${\lambda}>\|\operatorname{col}(\{{d}_{i}\}_{{i}\in\mathcal{I}_{a}})\|_{\infty}/$ $\max{}^{+}(\sigma({\mathbf{J}}_{0}))$, then the neutral opinion ${z}=0$ is an unstable equilibrium of the Neural NOD. Moreover, the opinion state departs ${z}=0$ (locally) at an exponential rate with an arbitrarily small bias ${b}$,
  • When bias term ${b}\neq 0$, the neutral opinion ${z}=0$ is not an equilibrium of the Neural NOD and the pitchfork unfolds, i.e., the model is ultrasensitive at ${z}=0$.
Proof:
Linearizing Neural NOD at neutral opinion ${z}=0$ gives $\dot{{z}}=(-D+{\mathbf{J}}_{0}){z}+{b}$, where $D=\operatorname{blkdiag}(\{D^{i}\}_{i\in\mathcal{I}_{a}})$. When ${b}=0$, indecision ${z}=0$ is an equilibrium. Since ${}^{+}(\sigma({\mathbf{J}}_{0}))\neq\emptyset$, if the attention variable ${\lambda}$ is chosen such that ${\lambda}>{\lambda}^{*}=\|\operatorname{col}(\{{d}_{i}\}_{{i}\in\mathcal{I}_{a}})\|_{\infty}/\max{}^{+}(\sigma({\mathbf{J}}_{0}))$, then matrix $-D+{\mathbf{J}}_{0}$ has one eigenvalue with positive real part, i.e., indecision ${z}=0$ is exponentially unstable. Thus, an indecision-breaking pitchfork bifurcation occurs at critical attention value ${\lambda}^{*}$. When ${b}\neq 0$, the pitchfork bifurcation unfolds, as predicted by the unfolding theory [golubitsky1985singularities]. ∎
IV Simulation Results
We use our proposed Neural NOD model, trained on both synthetic and human datasets, to automatically tune the game cost weights within the ILQGame planning framework [fridovich2020efficient]. We evaluate its performance in simulated autonomous racing scenarios (our running example). For both planning and simulation, we employ the 4D kinematic bicycle model [zhang2020optimization], discretized with a time step of $\Delta t=0.1$ s, to describe the vehicle motion. We implement the inverse game training pipeline with Flax [flax2020github] and train all neural networks with Adam [kingma2014adam]. All planners are implemented using JAX [jax2018github] and run in real-time at a frequency of 10Hz on a desktop with an AMD Ryzen 9 7950X CPU.
Hypotheses. We make three hypotheses that showcase the strengths of Neural NOD.
  • H1 (Performance). The Neural NOD model leads to safer and more competitive robot motion.
  • H2 (Generalization). The Neural NOD model generalizes better to out-of-distribution rival behaviors.